You have 235.5 g of Ba(OH)2.
Part 1
How many moles of barium hydroxide, Ba(OH)2, are present in 235.5 g?

mol Ba(OH)2
Part 2
How many moles of barium ions are present in Part 1?

mol Ba2+
Part 3
How many moles of hydroxide ions are present in Part 1?

mol OH1-
1
Expert's answer
2020-10-12T13:33:13-0400

PART 1

Ba = 137.327g


O = 16g


H = 1.008g


Molar mass Ba(OH)2 = (137.327+2(16)+2(1.008))g


Ba(OH)2 = 171.34g


Therefore : 1 Mol = 171.343g

? = 235.5g


= (235.5/171.343) = 1.4Moles


PART 2


1 mol of Ba+ = 137.327

? = 171.343

=1.2moles


PART 3

1.4 moles = 2mol


? = 1 mol


(1.4/2) =0.7 moles